QUESTION 4 3 charges are positioned along a horizontal line. A +5 C charge is placed at the point X 0 m A +7 C charge is placed at the point X : 0.4 m A-3 C charge is placed at the point X 1.3 m. What is the magnitude of the force on the +5 C charge, in units of Newtons?
Answer #1

force on +5C by+7C = kq1q2 / r1^2

= 9*10^9*5*7 / 0.4^2

= 1.968*10^12 N ( to the left)

force on +5C by -3 C = kq1q2/r^2

= 9*10^9*5*3 / 1.3^2

= 7.98*10^11 N ( to the right)

net force on charge +5C = (1.968*10^12 - 7.98*10^10) = 1.88*10^12 N ( to the left)